A study discovered that Americans consumed an average of 13.5 pounds of chocolate per year. Assume that the annual chocolateThe answers are in red. Please explain all parts, I really want to understand how to tell which quartile to use like part e!
Answer #1

a)
P( X<9) = ?
  
I know that, z = (X-mean)/(sd)  
z1 = (9-13.5)/3.9) = -1.1538
hence,  
P( X<9)   = P(Z<-1.1538) = NORMSDIST(-1.1538) = 0.1243

b)
P( X>11) = 1 - P(X<11) = ?
  
I know that, z = (X-mean)/(sd)  
z1 = (11-13.5)/3.9) =   -0.6410
  
hence,  
P( X>11) = 1- P(Z<-0.641) = 1 - NORMSDIST(-0.641) =   0.7392

c)
P(10 < X < 14) = P(X<14) - P(X<10) = ?
  
I know that, z = (X-mean)/(sd)  
z1 = (10-13.5)/3.9) =   -0.8974
z2 = (14-13.5)/3.9) =   0.1282
  
hence,  
P(10 < X < 14) = P(Z<0.1282) - P(Z<-0.8974) = NORMSDIST(0.1282) - NORMSDIST(-0.8974) = 0.3663

d)
P(X=12) = 0.0000
probability at a point is zero as normal is continuous variable

e)
P(Z<z) = 80%
z = NORMSINV(0.8)
z = 0.841621234
(X-mean)/sd =   0.8416
X = 0.8416*3.9+13.5
X = 16.8